OpenStudy (anonymous):

The city purchased 24 new buses that each hold 32 passengers. It also purchased 18 new smaller buses that each hold 27 passengers. What is the greatest number of passengers that could be on all the new buses if all the seats were filled

OpenStudy (mrnood):

1 bus with 32 passengers = 1* 32 = 32 people 2 bus with 32 passengers = 2* 32 = 64 people 3 bus with 32 passengers = 3* 32 = 96 people . . 24 buses with 32 passengers = ?* 32 = ?? people 1 bus with 27passengers = 1* 27= 27people 2 bus with 27passengers = 2* 27= 54people 3 bus with 27passengers = 3* 27= 81people . . 18 buses with 27passengers = ?* 27= ?? people Once you have the total for each type of bus - just add them up for the answer...

OpenStudy (anonymous):

486

OpenStudy (mrnood):

That is not the correct answer - if you show some working I may be able to help you see where you went wrong

OpenStudy (mrnood):

486 is the number that can be carried in the smaller buses. You need to add on the number from the larger buses

OpenStudy (anonymous):

so 24x32=768

OpenStudy (mrnood):

now add the 2 types of bus together

OpenStudy (anonymous):

so 768+486=1,254

OpenStudy (mrnood):

ok - do you understand how you did that?

OpenStudy (anonymous):

yes

OpenStudy (mrnood):

good - well done

OpenStudy (anonymous):

ok byee thanks for helping